a ball is kicked with an initial velocity of 11.1 m/s in the horizontal direction and 13.5 m/s in the vertical direction. At what speed does the ball hit the ground? Assume a level ground. What maximum height is attained by the ball?


1
Expert's answer
2021-09-04T07:34:52-0400

Given:

"v_x=11.1\:\rm m/s"

"v_y=13.5\:\rm m/s"


At the ground level the final speed is equal to initial speed

"v_f=v_i=\sqrt{v_x^2+v_y^2}=\sqrt{11.1^2+13.5^2}=17.5\:\rm m/s"

The maximum height is attained by the ball is given by

"h_{\max}=\frac{v_y^2}{2g}=\frac{13.5^2}{2*9.8}=9.30\:\rm m"